OpenStudy (anonymous):

how do I solve this system of linear equations {y=x+1 {y=2x-1 do I use in number ? how do you no which number to sub. for x and y please help

OpenStudy (anonymous):

ok, here y= either x+1, or 2x-1. therefore, you can set both equal to each other: 2x-1=x+1 x-1=1 (subtract x on both sides) x=2 (and 1 to both sides) you can check if this is correct by substituting 2 in both functions, and see if you get the same value

OpenStudy (anonymous):

add as a fan please :DD